### v1

@@ Abstract
Experiments deliver credible treatment effects estimates, but are often localized to specific sites, populations, or mechanisms. When such estimates are insufficient to extrapolate effects for broader policy questions as for external validity and general-equilibrium (GE) effects, researchers combine trials with external evidence from reduced-form or structural observational estimates, or prior experiments. We develop a unified framework for designing experiments in this setting: the researcher selects which parameters to identify experimentally from a feasible set (which treatment arms and/or individuals to include in the experiment), allocates sample size, and specifies how to weight experimental and observational estimators. Because observational inputs may be biased in ways unknown ex ante, we adopt a minimax proportional-regret objective that evaluates any candidate design relative to an oracle that knows the bias and jointly chooses the design and estimator. This yields a transparent bias-variance trade-off that requires no prespecified bias bound and depends only on information about the precision of the estimators and the estimand’s sensitivity to underlying parameters. We illustrate the framework by (i) designing small-scale cash-transfer experiments aimed at estimating GE effects and (ii) optimizing site selection for microfinance interventions.
@@ 1 Introduction
Randomized controlled trials have transformed empirical economics by delivering credible, internally valid estimates. However, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism. These effects, although useful, are often not sufficient to answer broader questions about external validity, generalizability, or equilibrium effects. In response, a growing literature in development economics (e.g., Meghir et al., 2022; Gechter et al., 2024; Bassi et al., 2022; Attanasio et al., 2012), political economy (e.g., de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), and labor economics (e.g., Chetty et al., 2016) complements localized experiments with external evidence from reduced-form or structural observational estimates, and/or trials in other contexts, with the goal of estimating complex counterfactuals that no single experiment can fully identify. For example, in our review of AEA journals (in 2015-2025), over 30% of experimental papers report, in addition to the experiment, at least one observational estimate – either reduced-form (e.g., via matching or instrumental variable regression), or structural (see Figure 12). This raises a design question: given constraints on what experiments can identify, which experiments should be run (and how) when they will be combined with external evidence to estimate those counterfactuals?
@@
To answer these questions, we develop a framework for designing experiments to be used alongside external evidence. By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other times or places. Our main contribution is a joint procedure that selects which experiments to run subject to a budget (i.e., which treatment arms and/or sub-populations to include in the study), sets their precision via sample allocation, and prescribes how to combine experimental estimates with external evidence. A central consideration in our design is that external inputs may be biased in ways not known ex ante.
We consider a setting where the object of interest is a known function $\tau(\theta)$ with $\theta$ a vector of unknown parameters; $\tau$ may be univariate or multivariate and typically encodes a policy-relevant target. Researchers have access to external estimates of $\theta$ that may be biased—henceforth, observational estimates. They then run one or more experiments in the population of interest to learn a subset of parameters without bias, choosing a sample allocation across sites and arms subject to a budget. We parameterize the design so that some, but not necessarily all, components of $\theta$ can be learned experimentally. After the experiment, each experimental estimate is combined with its observational counterpart via parameter-specific shrinkage weights (chosen jointly with the design), while parameters not learned experimentally rely on the observational estimates alone.
@@
Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in a region receive the subsidy. The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium, and (iii) wage/price adjustments that shift the returns to schooling. A large-scale trial that identifies GE effects is often infeasible; instead, one can run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$. For instance, researchers may choose between two (small-scale) treatment arms such as a CCT to estimate the direct effect or an unconditional-cash arm to measure the income elasticity, and extrapolate the GE effects using price effects from other experimental or observational studies. When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ may target the average effect across sites.
@@
A natural starting point is to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps in a worst-case sense over bias in the observational estimates. In practice, however, the size and even the direction of bias are unknown (before the experiment is run). A worst-case approach would therefore be overly-conservative and disregard information about the variance. We instead adapt the definition of proportional (adaptation) regret previously studied when combining estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem. Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the (shrinkage) estimator. Taking the supremum of this ratio over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
@@
Our main result gives an explicit, easy-to-compute characterization of proportional regret. For any design and shrinkage weights, the regret is the maximum of two normalized components. The first is a variance component: the estimator’s sampling variance under the chosen design and weights, divided by the smallest variance attainable by any feasible design. The second is a bias component: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias. The variance component depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters. This gradient is evaluated, under mild conditions, at the observational estimates. [^4] For this result to hold, we assume that $\tau(\theta)$ is a smooth and differentiable function in $\theta$. For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al. (2020); Bonhomme and Weidner (2022), but consider a more general framework where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error. Section 4.1 provides details. By contrast, the bias component depends only on the sensitivity vector and the shrinkage weights, both known ex-ante.
@@
This characterization makes the bias-variance trade-off transparent. At the optimum, the two normalized components tend to be equalized: when the bias component dominates, weight shifts toward experimental evidence and the design prioritizes learning high-sensitivity coordinates; when variance dominates, the procedure invests sample where it most reduces variance and relies more (via shrinkage) on the most precise observational inputs. This yields a nested procedure: we (i) find (regret-optimal) shrinkage weights for each candidate design to balance the two ratios; (ii) allocate sample, subject to the budget, to reduce the variance; and (iii) select the experiment set, trading off designs that improve precision the most against those targeting parameters to which the policy estimand is the most sensitive. The resulting design and estimator only require knowledge of the expected variance-covariance matrix (standard in experimental design problems, Gerber and Green, 2012) and of the observational estimates. It can then be reported in a pre-analysis plan.
@@
We illustrate the framework in the cash‐transfer application. As an illustrative example, we consider a researcher evaluating a conditional cash transfer (CCT) to increase schooling in rural Kenya. Budget and feasibility constraints permit only small-scale (partial-equilibrium) randomization, while the goal is to predict general-equilibrium effects. As preliminary (external) inputs, the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006). We are concerned that these preliminary estimates may lack external validity in Kenya. We compare three designs: (i) a CCT arm to estimate the direct schooling effect; (ii) an unconditional cash transfer (UCT) arm to estimate income effects; and (iii) a two-arm design that allocates sample across CCT and UCT under a common budget. Running a single arm captures settings with fixed costs per arm; allowing both arms with budgeted allocation captures cases with non-binding fixed costs but binding variable costs. When both arms are feasible, the regret-optimal allocation puts most participants in the CCT and a small fraction in the UCT, reflecting the target’s greater sensitivity to the direct effect despite the CCT’s higher variance. If only one arm can be run, the choice hinges on sample size: for very small $n$, the lower-variance of the UCT arm makes this preferred over a (too noisy) CCT arm; once $n\geq 500$, the more informative but noisier CCT yields lower regret. Relative to an oracle that knows the bias, the optimal design closely tracks oracle performance, and regret approaches one as sample size grows.
@@
In summary, this paper links the experimental-design literature—which typically focuses on settings where all parameters are identified within the experiment and leaves aside questions of data combination—to recent work that integrates experimental and (reduced-form or structural) observational evidence to extrapolate effects in complex scenarios.
@@
Recent advances for experimental design include balancing and variance-minimizing allocations (Tabord-Meehan, 2018; Bai, 2019; Cytrynbaum, 2021; Kallus, 2018; Bertsimas et al., 2015), adaptive designs for policy choice (Kasy and Sautmann, 2019; Russo et al., 2018; Kato et al., 2024; Cesa-Bianchi et al., 2025), and experimental design under correct model specification (Silvey, 2013; Chaudhuri and Mykland, 1993; Chaloner and Verdinelli, 1995; Higbee, 2024; Kiefer and Wolfowitz, 1959; Reeves et al., 2024; Viviano, 2020; Kasy, 2016). Traditional work on experimental design for robust-model estimation either focused on testing competing models (Atkinson and Fedorov, 1975; López-Fidalgo et al., 2007), or on using a-priori knowledge of (worst-case) bias for the design of an experiment (Box and Draper, 1959; Wiens, 1998; Tsirpitzi et al., 2023; Sacks and Ylvisaker, 1984). We complement this literature by introducing and studying the question of which experiment to design when combined with observational evidence (when not all parameters can be learned experimentally).
@@
Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design. References include Manski and Tetenov (2016), Banerjee et al. (2020), Manski (2004), Dominitz and F. Manski (2017), Olea et al. (2024), Hu et al. (2024), Breza et al. (2025) among others. These focus on settings where researchers have only access to experimental variation. Here instead we study the problem of choosing which experiment to run (and how) when combined with observational evidence. We also complement literature that leverages correctly-specified models for site selection (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
@@
A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2017, 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024). We adapt the worst-case proportional regret idea from Armstrong et al. (2024) and Tsybakov (1998) here studied in the context of experimental design. Our definition of sensitivity of the estimand to each parameter directly links to sensitivity analysis in Andrews et al. (2020). Unlike both strands of this literature where the design is fixed and researchers optimize over the choice of the estimator only, here we optimize the design itself. Consequently, the objective changes: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
Finally, we connect to a growing empirical literature that combines observational (or model-based) evidence with experimental variation to learn economic quantities – spanning general equilibrium effects (Egger et al., 2022; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023), external-validity counterfactuals (Gechter et al., 2024; Banerjee et al., 2024), large-scale information campaigns (de Albuquerque et al., 2025; Larroucau et al., 2024), and market-system experiments (Bergquist and Dinerstein, 2020; Allende et al., 2019). We provide a unifying framework to decide what to learn experimentally, how to split the sample across experiments, and how to integrate experimental and observational evidence.
@@ 2 Problem description
Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$. The estimand may arise from a reduced-form (e.g., treatment effect in a given area) or a structural model. For simplicity, we assume that the estimand is univariate, while Section 4.3 shows how our framework directly extends to multivariate $\tau(\theta)\in\mathbb{R}^{q}$ for $q>1$.
@@
2.1 Setup
We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$. We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model. However, $\tilde{\theta}^{\mathrm{obs}}$, that for exposition we define observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$. Examples of $\tilde{\theta}^{\mathrm{obs}}$ include a structural model estimate with confounding, an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity.
Researchers may also collect evidence for a subset of parameters that they know is unbiased. For instance, researchers may generate exogenous variation via an experiment (or more generally acquiring more data) that identifies one or some of the parameters of interest. We refer to these parameters as experimental estimates, adopting here the definition of experimental estimates as parameters that the researcher knows are unbiased. [^5] In some applications researchers may be confident that a few observational estimates have no bias. It is possible to extend our framework to these settings with a change of notation where those observational estimates assumed to be unbiased are defined as experimental estimates, and such experimental estimates are always contained in the set $\mathcal{E}$ defined below.
Setting 1 (Main setup).
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a known strictly positive-definite matrix $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})\subset\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$ with uniformly bounded entries, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
@@ 2.1 Setup
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance as
$\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E}).$ (1)
Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators
@@
Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design (which parameters are learned and with what precision). When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$.

### v2

@@ Abstract
Experiments deliver credible treatment effects estimates, but are often localized to specific sites, populations, or mechanisms. When such estimates are insufficient to extrapolate effects for broader policy questions as for external validity and general-equilibrium (GE) effects, researchers combine trials with external evidence from reduced-form or structural observational estimates, or prior experiments. We develop a unified framework for designing experiments in this setting: the researcher selects which parameters (or moments) to identify experimentally from a feasible set of possible experiments, allocates sample size, and specifies how to weight experimental and observational estimators. Because observational inputs may be biased in ways unknown ex ante, we adopt a minimax proportional-regret objective that evaluates any candidate design relative to an oracle that knows the bias and jointly chooses the design and estimator. This yields a transparent bias-variance trade-off that requires no prespecified bias bound and depends only on information about the precision of the estimators and the estimand’s sensitivity to underlying parameters. We illustrate the framework by (i) designing small-scale cash-transfer experiments aimed at estimating GE effects and (ii) optimizing site selection for microfinance interventions.
@@ 1 Introduction
Randomized controlled trials (RCTs) have improved empirical economics by providing internally valid causal estimates. However, feasibility constraints often confine trials to localized effects such as the effect in a specific site or subpopulations, or of a particular mechanism. These effects, although useful, are often not sufficient to answer broader questions about external validity and (general) equilibrium effects of at-scale interventions. For instance, the median evaluation of experiments in top economics journals is “representative of a population of 10,885 units, studied a treatment delivered to 5,340 units, and randomized in clusters of 26 units per cluster” (Muralidharan and Niehaus, 2017).
In response, a growing literature in development economics (e.g., Meghir et al., 2022; Bassi et al., 2022; Attanasio et al., 2012; de Albuquerque et al., 2025), education (e.g., Allende et al., 2019; Larroucau et al., 2024), labor economics (e.g., Chetty et al., 2016), and industrial organization (e.g., Katz and Allcott, 2025; Allcott et al., 2025) combine experiments with external evidence—observational estimates and/or results from experiments in other contexts—to estimate complex counterfactuals that no single experiment can identify. In our review of AEA journals over the past decade, more than 30% of experimental papers complement RCT evidence with at least one observational input, either reduced-form or structural (Figure 12). This trend poses a design question: given constraints on feasible experiments, which experiments should be run (and how) to estimate such counterfactuals?
Our main contribution is a framework for designing experiments combined with external evidence. By external evidence we mean reduced-form or structural estimates based on observational data, and/or results from experiments conducted in other contexts. We provide a procedure that selects which experiments to run subject to a budget (e.g., which treatment arms and/or sub-populations to include), sets their precision via sample allocation, and prescribes how to combine experimental and external evidence. External input may be biased in ways not known ex ante (e.g., due to confounding or lack of external validity).
@@
The first step is to specify the estimand and the underlying parameters. Define $\tau(\theta)$ the (policy-relevant) object of interest, where $\tau$ is a known smooth function and $\theta$ is a vector of unknown parameters; $\tau$ may be scalar or vector-valued. Researchers also observe external baseline estimates of $\theta$—henceforth observational estimates—which may be biased. Each feasible experiment can identify a subset of parameters without bias. This parameterization makes explicit which components of $\theta$ are learned experimentally and which may rely on biased observational evidence; not all components need (or can) be learned experimentally.
@@
Returning to the cash-transfer example, $\tau(\theta)$ is the effect on schooling when all poor households in rural Kenya receive the subsidy. The parameter vector $\theta$ bundles (i) the direct schooling response to a conditional cash transfer (CCT), (ii) the income effect of transfers in partial equilibrium , and (iii) wage adjustments that shift the returns to schooling. Due to cost constraints, researchers may only run small, partial-equilibrium experiments and then combine these with prior evidence to recover $\tau(\theta)$. They may choose between a CCT treatment arm to estimate the direct effect, or an unconditional-cash arm (UCT) to measure the income elasticity, and extrapolate the GE effects using evidence from a previous study in Mexico (Todd and Wolpin, 2006). (When site selection is also part of the design, $\theta$ additionally collects site-specific effects and $\tau(\theta)$ targets the average effect across sites.)
The second step is to design the experiment: (i) selecting which experiment to run; (ii) allocating sample sizes across arms; and (iii) choosing how to combine observational and experimental evidence. We allow flexibility in (iii): researchers may choose to either optimally combine both sources when available or rely solely on experimental estimates. Parameters not identified experimentally are based on observational estimates.
@@
A natural starting point is to design the experiment to minimize the mean-squared error (MSE) for estimating $\tau(\theta)$, perhaps worst-case over the observational bias. In practice, however, the size and even the direction of bias are unknown (before the experiment). A worst-case approach would therefore be overly-conservative and disregard information about the variance. We instead adapt the definition of proportional (adaptation) regret previously studied when combining two estimators generated by a fixed design (Armstrong et al., 2024; Tsybakov, 1998) and generalize it to the experimental design problem (with multiple estimators). Here, the regret is the ratio of the worst-case researcher’s mean-squared error relative to an oracle that knows the worst-case bound on the observational bias and chooses both the design and the estimator. Taking its supremum over any value of the bias yields a procedure that is robust to any level of misspecification without requiring priors about the bias.
@@
Our main result gives an explicit novel characterization of proportional regret. For any design and weights that combine the experimental and observational point estimates, the regret is the maximum of two normalized components. The first is a variance regret: the estimator’s sampling variance under the chosen design and weights, divided by the smallest attainable variance. The second is a bias regret: the worst-case squared bias induced by using external evidence, divided by the smallest feasible worst-case bias over the class of designs. The variance regret depends on the (expected) variance-covariance matrix of the observational and experimental estimates implied by the design, the shrinkage weights, and the sensitivity of the policy parameter, i.e., the gradient of $\tau$ with respect to the parameters. This gradient is evaluated, under mild conditions, at the observational estimates. [^4] For this result to hold, we assume that $\tau(\theta)$ is twice differentiable with bounded derivatives. For non-linear $\tau(\theta)$ in $\theta$, we also require that the observational estimates bias is local to zero in the spirit of Andrews et al. (2020); Armstrong and Kolesár (2021); Bonhomme and Weidner (2022), where its rate of convergence can be the same, faster, or even slower than the estimators’ standard error. Section 5.1 provides details. The bias regret depends only on the sensitivity vector and the shrinkage weights, known ex-ante.
@@
This characterization makes the bias-variance trade-off transparent. At the optimum, the two normalized components tend to be equalized: when the bias component dominates, the design prioritizes high-sensitivity parameters; when variance dominates, the procedure invests sample where it most reduces variance and relies more on the most precise observational inputs. This yields a nested procedure that chooses jointly the estimator, the sample allocation and the experiment to run, trading off precision with sensitivity of the policy estimand. The resulting design and estimator only require knowledge of the expected variance-covariance matrix (standard in experimental design problems, Gerber and Green, 2012) and of the observational estimates. It can then be reported in a pre-analysis plan. That is, the inputs for the design are the same that would be required by standard power-analysis plans.
From a theoretical perspective, our result leverages a key insight. Because here bias arise from multiple parameters, the choice of the uncertainty set is pivotal. While we do not require researchers to specify such a set, regret is with respect to an oracle that knows the worst-case bias within an $\ell_{\infty}$-ball. Compared with $\ell_{1}$ or $\ell_{2}$, the $\ell_{\infty}$-ball prevents that a large bias in one component is “offset” by smaller biases elsewhere. This is attractive when biases may be positively correlated across parameters. We show that the $\ell_{\infty}$ choice combined with our design problem (with decisions made before experimental data is realized), yields a quasi-convex regret in the bias radius as the oracle optimizes over the designs. This result, worst-case over the radius, delivers a transparent and easy-to-compute variance-bias trade-off.
We then extend the model to general method-of-moments estimators that combine experimentally identified moments with observational moments. As in the baseline case, observational moments may incur unknown misspecification. We optimize (i) which moments to identify experimentally and (ii) how to combine experimental and observational moments to estimate the counterfactual of interest. Our main result continues to hold in this more general scenario provided that the moment estimators are (asymptotically) linear in the moments. Linearity occurs (under smoothness conditions) as long as the ratio of the bias to standard error is no faster than $n^{1/4}$, with $n$ denoting the observational sample size; standard local-asymptotics with the bias and standard error of the same order is a special case.
@@
We illustrate the framework in the cash‐transfer application, when the researcher uses evidence from Mexico’s PROGRESA program (Todd and Wolpin, 2006; Attanasio et al., 2012). We are concerned that these preliminary estimates may lack external validity in Kenya. We compare three designs: (i) a CCT arm to estimate the direct schooling effect; (ii) an unconditional cash transfer (UCT) arm to estimate income effects; and (iii) a two-arm design that allocates sample across CCT and UCT under a common budget. Running a single arm captures settings with fixed costs per arm; allowing both arms with budgeted allocation captures cases with non-binding fixed costs but binding variable costs. When both arms are feasible, the regret-optimal allocation puts most participants in the CCT and a small fraction in the UCT, reflecting the target’s greater sensitivity to the direct effect despite the CCT’s higher variance. If only one arm can be run, the choice hinges on sample size: for very small $n$, the lower-variance of the UCT arm makes this preferred over a (too noisy) CCT arm; once $n\geq 500$, the more informative but noisier CCT yields lower regret. Relative to an oracle that knows the bias, the optimal design closely tracks the oracle. We provides a step-by-step guide for practitioners applicable in other contexts.
@@
This paper links the experimental-design literature—which mostly focuses on settings where all parameters are identified within the experiment and leaves aside questions of data combination—to recent work that integrates experimental and (reduced-form or structural) observational evidence to extrapolate effects in complex scenarios.
@@
Recent advances for experimental design include balancing and variance-minimizing allocations (Tabord-Meehan, 2018; Bai, 2019; Cytrynbaum, 2021; Kallus, 2018; Bertsimas et al., 2015), adaptive designs for policy choice (Kasy and Sautmann, 2019; Russo et al., 2018; Cesa-Bianchi et al., 2025), and experimental design under correct model specification (Silvey, 2013; Chaudhuri and Mykland, 1993; Chaloner and Verdinelli, 1995; Higbee, 2024; Kiefer and Wolfowitz, 1959; Reeves et al., 2024; Viviano, 2020; Kasy, 2016). Traditional work on experimental design for robust-model estimation either focused on testing competing models (Atkinson and Fedorov, 1975; López-Fidalgo et al., 2007), or on using a-priori knowledge of (worst-case) bias for the design of an experiment (Box and Draper, 1959; Wiens, 1998; Tsirpitzi et al., 2023; Sacks and Ylvisaker, 1984). All these references leave aside questions about observational data combination. In the context of minimizing the variance of an experiment, Rosenman and Owen (2021) proposes to use observational studies to construct high-confidence bounds on the experimental variance, and then focus on variance-optimal stratification. This problem differs from our design problem and analysis, which instead studies the use of observational studies in combination with experiments for extrapolation.
In summary, we complement this literature by studying which experiment to run (and how) when not all parameters relevant for the target estimand can be learned experimentally.
@@
Our minimax regret criterion connects to a long-standing decision-theoretic tradition for experimental design. References include Manski and Tetenov (2016), Banerjee et al. (2020), Manski (2004), Dominitz and F. Manski (2017), Olea et al. (2024), Hu et al. (2024), Breza et al. (2025) among others. These references focus on settings where researchers have only access to experimental variation, instead of combining it with observational evidence, motivating different designs and objective functions. In the context of site-selection, we complement literature that leverages correctly-specified models (Gechter et al., 2024; Abadie and Zhao, 2021) by allowing for misspecification in observational estimates.
@@
A related line of work analyzes estimation under misspecification (Armstrong and Kolesár, 2018; Armstrong et al., 2024; Andrews et al., 2025; Andrews et al., 2020; Bonhomme and Weidner, 2022; Christensen and Connault, 2023; James et al., 1961), and combining existing experiments with observational studies (Gechter, 2022; Athey et al., 2025, 2020; Kallus et al., 2018; Dutz et al., 2021; Rosenman et al., 2022; Bhattacharya, 2013; de Chaisemartin and D’Haultfœuille, 2020; Rambachan et al., 2024). We adapt the worst-case proportional regret idea from Armstrong et al. (2024) and Tsybakov (1998) here studied in the context of experimental design. Our definition of sensitivity of the estimand to each parameter builds on sensitivity analysis in Andrews et al. (2020). Unlike both strands of this literature where the design is fixed and researchers optimize over the choice of the estimator only, here we optimize the design itself. The objective and analysis change: regret is computed against the best design-estimator combination, and the resulting estimator is learned jointly with the design.
@@ 2 Problem description
Consider a researcher interested in an arbitrary target estimand $\tau(\theta)\in\mathbb{R}$, indexed by a low-dimensional unknown parameter vector $\theta\in\mathbb{R}^{p}$ and a known mapping $\tau$. For simplicity, we assume that the estimand is univariate, while Section 5.3 shows how our framework directly extends to multivariate $\tau(\theta)\in\mathbb{R}^{q}$ for $q>1$.
@@
Basic setup
We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$. We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model. However, $\tilde{\theta}^{\mathrm{obs}}$, defined as observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$. Examples include an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity. Researchers can also collect experimental evidence for a subset of parameters, producing estimates whose bias equals zero by design.
Setting 1.
For a subset $\mathcal{E}\subseteq\{1,\ldots,p\}$ and a known positive-definite matrix $\Sigma(\mathcal{E})\in\mathbb{R}^{(p+|\mathcal{E}|)\times(p+|\mathcal{E}|)}$, define $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$ an observational estimate and $\tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}\in\mathbb{R}^{|\mathcal{E}|}$ an experimental estimate each satisfying
@@
for an unknown bias vector $b\in\mathbb{R}^{p}$. Moreover, define its joint variance-covariance matrix as $\mathbb{V}\!\begin{pmatrix}\tilde{\theta}^{\mathrm{obs}}-\theta\\[2.0pt] \tilde{\theta}^{\mathrm{exp}}_{\mathcal{E},\Sigma}-\theta_{\mathcal{E}}\end{pmatrix}=\Sigma(\mathcal{E}).$ Given $(\mathcal{E},\Sigma)$, consider a class of linear plug-in estimators
@@
Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design (which parameters are learned and with what precision). When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$. We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012) (in practice, any consistent estimator, e.g., from pilot studies would suffice).
The class of experiments and estimators is assumed to be in a user-specific set $\mathcal{D}$.
Assumption 1 (Feasible experiments and estimators).
We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible design (i.e., choice of $\mathcal{E}$ and $\Sigma(\mathcal{E})$) and estimators (i.e., $\gamma$). The set $\mathcal{D}$ can be arbitrary as long as (i) each feasible $\Sigma$ has uniformly bounded entries and is strictly positive definite; (ii) $\gamma_{j}=1$ for all $j\in\mathcal{E}$ is one admissible choice (among potentially other values).